On April 1, 2017, Teal Company sold 13,500 of its 11%, 15-year, $1,000 face value bonds at 97. Interest payment dates are April 1 and October 1, and the company uses the straight-line method of bond discount amortization. On March 1, 2018, Teal took advantage of favorable prices of its stock to extinguish 6,300 of the bonds by issuing 207,900 shares of its $10 par value common stock. At this time, the accrued interest was paid in cash. The company’s stock was selling for $31 per share on March 1, 2018.

Prepare the journal entries needed on the books of Teal Company to record the following. (Round intermediate calculations to 6 decimal places, e.g. 1.251247 and final answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

(a) April 1, 2017: issuance of the bonds.
(b) October 1, 2017: payment of semiannual interest.
(c) December 31, 2017: accrual of interest expense.
(d) March 1, 2018: extinguishment of 6,300 bonds. (No reversing entries made.)

No.

Date

Account Titles and Explanation

Debit

Credit

(a)

4/1/17

(b)

10/1/17

(c)

12/31/17

(d)

3/1/18

(To record payment to retiring bondholders)

3/1/18

(To record extinguishment of the bonds)
Answer #1

No.

Date Account Titles and Explanation Debit Credit
(a) 42826 Cash (13500*1000*97%) 13095000
Discount on bonds payable 405000
Bonds Payable (13500*1000) 13500000
(b) 43009 Interest Expense (13500+742500) 742500
Discount on bonds payable (405000/(15 years * 2)) 13500
Cash (13500,000*11%*1/2) 742500
(c) 43100 Interest Expense 378000
Discount on bonds payable (13500/2) 6750
Interest Payable (13500,000*11%*3/12) 371250
(d) 43160 Interest Payable (371250/13500*6300) 173250
Interest Expense 117600
Discount on bonds payable ((6750/3*2)/13500*6300) 2100
Cash 288750
((742500/6 months*5 months)/13500 bonds *6300)
(To record payment to retiring bondholders)
(d) 43160 Bonds Payable (6300*$1000) 6300000
Loss on bond redemption 322350
Discount on bonds payable (Working) 177450
Common Stock (207900 shares x $10) 2079000
Additional paid in capital (207900 shares x( $31-10)) 4365900
(To record extinguishment of the bonds)

Working:

Discount on Bonds Payable =
Discount on 6300 bonds = =405000/13500*6300 = $189000
Discount for 1 month = $189000 / 30 = $1050 per month
Unamortized discount for 169 months (180-11) months = $1050*169 = $177450